Gas Laws:
Part A
What will the pressure inside the container become if the piston is moved to the 1.60 L mark while the temperature of the gas is kept constant?
Express your answer with the appropriate units.

Part B
The gas sample has now returned to its original state of 1.00 atm, 20.0 ∘C and 1.00 L. What will the pressure become if the temperature of the gas is raised to 200.0 ∘C and the piston is not allowed to move?
Express your answer with the appropriate units.
Part C
The gas described in parts A and B has a mass of 1.66 grams. The sample is most likely which monoatomic gas?
Type the elemental symbol of the gas below.
1
Expert's answer
2017-03-20T10:01:06-0400
A) What was the starting pressure? You can calculate the final pressure as P2 = P1*T2/T1

B) The isochoric process: P2 = P1*T2/T1 = 1atm*473K/293K = 1.61 atm.

C) Molar mass = mRT/PV = 1.66g*8.314J/mol*K293K/101.3kPa*0.001m3 = 39.8 g/mol or Ar(argon)